OpenStudy (anonymous):

1,2,4,7,11 rule, next three numbers

myininaya (myininaya):

1+1=2 2+2+4 4+3=7 7+4=11 11+5=16 16+6=22 22+7=29

OpenStudy (anonymous):

but it needs a rule, so what would the rule be?

myininaya (myininaya):

f_n=f_(n-1)+n where f_1=2 and f_2=f_(2-1)+2=f_(1)+2=2+2=4 f_3=f_(3-1)+3=f_(2)+3=4+3=7 ok so f_n=f_(n-1)+n is a recursive pattern

myininaya (myininaya):

for n>=2

myininaya (myininaya):

and n is an integer

OpenStudy (anonymous):

none of this seems right to me, can you put it into y's and x's?

OpenStudy (anonymous):

1/2(n^2-n+2) n are natural numbers

myininaya (myininaya):

la-mrc that also looks good